Two individuals are involve in a collective business, each year Mr.A received 60 % of the profit while B receive 35% of remaining. The leftover profit is allocate for charity. The annual revenue of the business is $500,000 while the operating cost is $-175000 for inputs (including labor and capital), $-70,000 for rent, and 15% of rent as utilities. By profession Mr. A is an engineer and can earn $7,000 per month by working in automobile industry. Similarly, Mr.B can earn $5000 per annum by working in a nearby school. Assuming Zero income and sale tax, answer following questions.

a) What would be business and economic profit of Mr. A? b) Which shareholder is economically well-off and Why? c) What would be their total revenue in 15 years if the growth rate of the business is 2.3% per annum?
1
Expert's answer
2017-09-25T09:40:07-0400
Answer
• Business profit=(500000-175000- 70000-(70000*0.15))*0.6=146700
Economic profit=146700-7000*12=62700
• Shareholder Mr. A, because his economic profit more than economic Mr. B profit.
• 500000*(1+0.023)^15=703241.53
Question 2
The total business value of KFC in 1990 was $ 106,940. The management is expecting
growth of 1.3 percent.
How long KFC business would take to reach at $154,287 worth
Answer:
106940*(1+0.013)^x=154287
1.013^x=1.4427
X=28.37